OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. R(0, _____ ), S(-4, 2) m = 5

OpenStudy (callisto):

\[slope =\frac{y_2-y_1}{x_2-x_1}= \frac{y-2}{0-(-4)} = 5\]\[\frac{y-2}{4} = 5\]Can you solve it?

OpenStudy (anonymous):

thank youu!!! really appreciate the explanation!!! <3

OpenStudy (callisto):

welcome :)

OpenStudy (anonymous):

wait im sure its not 22.....

OpenStudy (callisto):

m=5 is the slope, right?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Then.. it should be correct :|

OpenStudy (anonymous):

ohh ok:)

OpenStudy (anonymous):

T(13, 2), W( ___ , -5) m = 7/2 <----what would i do here???

OpenStudy (callisto):

\[slope =\frac{y_2-y_1}{x_2-x_1}= \frac{2-(-5)}{13-x} = \frac{7}{2}\]

OpenStudy (anonymous):

how would i do the next step? 3/ 13-x =7/2

OpenStudy (callisto):

\[\frac{2-(-5)}{13-x} = \frac{7}{2}\]\[\frac{2+5}{13-x} = \frac{7}{2}\]Cross multiplication\[2(2+5) = 7(13-x)\]

OpenStudy (anonymous):

16= 7(13-X)

OpenStudy (anonymous):

so thenn... i divide by 7?

OpenStudy (callisto):

2(2+5) = 7(13-x) 2(7) = 7 (13-x) Divide both sides by 7

OpenStudy (anonymous):

omg.. my math is not good today

OpenStudy (anonymous):

its 6 right

OpenStudy (anonymous):

dont tell me its not 6..... D:

OpenStudy (callisto):

It's not....

OpenStudy (callisto):

2(7) = 7 (13-x) 2 = 13 -x x = ... ?

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

11...

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

id give you another medal.. cause im soo not smart-.-

OpenStudy (callisto):

Nope :| Just be careful with the calculations !!

OpenStudy (anonymous):

okay:D